OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair. (x'y) 5x-3y=29 x-4y=16

OpenStudy (anonymous):

5x - 3y = 29 and x - 4y = 16 x = 16 + 4y 5(16+4y) - 3y = 29 80+20y - 3y = 29 80+17y = 29 17y = -51 y = -3 5x - 3(-3) = 29 5x + 9 = 29 5x = 20 x = 4